OpenStudy (anonymous):

Algebra 2! Quadratic equations that can be solved using the quadratic formula can also be solved by factoring. A. Always B. Never C. Sometimes

OpenStudy (anonymous):

C

OpenStudy (saifoo.khan):

C

OpenStudy (saifoo.khan):

Lol. wrong PAste

OpenStudy (anonymous):

i know the answer they are looking for is sometimes <.< but the true answer is always. But put sometimes.

OpenStudy (anonymous):

Joe is right if u use the quadratic equation technically u found the roots so u can use it to factor the equation. but yea put sometimes

jimthompson5910 (jim_thompson5910):

The answer is C) Sometimes Consider the quadratic equation x^2+5x+6=0, we can factor the left side into (x+2)(x+3). However, the quadratic equation x^2+3x+7=0 cannot be factored (over the rationals), but it can be solved using the quadratic formula.

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

true but the problem ddnt say over rationals so u could argue this 1 is what we were saying

jimthompson5910 (jim_thompson5910):

well my experience is when they say "factor" they mean "over the rationals" (at least with high school that is). Ie, find two rational numbers p and q such that ax^2+bx+c=(x-p)(x-q), but you have a point jahtoday.

myininaya (myininaya):

you should tell your teacher this is a trick question i think she would mean factor over rationals but it doesn't say that

myininaya (myininaya):

anything can be factored but it would be hard to get some of those factors that factor over the real/complex without using quadratic but i think it is possible with enough trail and error lol

myininaya (myininaya):

you know without using the quad formula

OpenStudy (anonymous):

if someone asked me what x^2+3x+7 factored into, i would answer: \[(x-\frac{-3-i \sqrt{19}}{2})(x-\frac{-3+i \sqrt{19}}{2})\] >.>

myininaya (myininaya):

lol

myininaya (myininaya):

see its easy lol

myininaya (myininaya):

yes thats what i would say to joe without using quadratic formula its an easy guess right saifoo?

OpenStudy (saifoo.khan):

Myinin, v both are going with same Pace, i just checkd. Same Level. Same Medals! :D

myininaya (myininaya):

lol

OpenStudy (saifoo.khan):

Oh Yea it is! :D

OpenStudy (saifoo.khan):

hehe.